Bath 1: 6’ by 10’
Bath 2: 8’ by 14’
Kitchen: 12’ by 14’
Find the areas of each of these, and then find the cost by multiplying the areas by $5/sq. ft.
can someone help me with this please

Answer by nerdybill(7384)

Bath 1: 6’ by 10’
area: 6*10 = 60 sq ft
cost: 60*5 = $300
.
Bath 2: 8’ by 14’
area: 8*14 = 112 sq ft
cost: 112*5 = $560
.
Kitchen: 12’ by 14’
area: 12*14 = 168 sq ft
cost: 168*5 = $840
.